OpenStudy (anonymous):

Which value of k makes (2x + 3)(kx – 5) = 2kx^2 – x – 15 true? A. –3 B. –1 C. 1 D. 3

OpenStudy (anonymous):

Please anyone help!!! :(

OpenStudy (johnweldon1993):

foil out those parenthesis

jimthompson5910 (jim_thompson5910):

Hint: (2x + 3)(kx – 5) = 2kx^2 – x – 15 2x(kx – 5) + 3(kx – 5) = 2kx^2 – x – 15 2kx^2 – 10x + 3kx – 15 = 2kx^2 – x – 15 -10x + 3kx – 15 = -x – 15 -10x + 3kx = -x I'll let you finish.

OpenStudy (anonymous):

I have no idea how to finish it :(

jimthompson5910 (jim_thompson5910):

you want to solve for k, so you need to isolate it

OpenStudy (johnweldon1993):

well...we see that our x's can combine no? -10x + 3kx = -x combine the x's add 10x to both sides and get 3kx = 9x now can you solve for k?

OpenStudy (anonymous):

k is equal to three?

OpenStudy (johnweldon1993):

very good :)

OpenStudy (anonymous):

THANK you so much :):):))))

OpenStudy (johnweldon1993):

no problem...do you understand it?

OpenStudy (anonymous):

After you kind of simplifying it I understand your steps and how you came to finding k, but I wouldn't have a clue how to start if you didn't show how